in a regression relationship between X and Y, if a = 2 and b = 4, what would the value of Y be when X = 1?
Answer #1

The regression equation is,

Y = a + b X

For, a = 2 and b = 4

Y = 2 + 4X

When X = 1,

Y = 2 + 4 * 1 = 6